Manny, Moe and Jack are among 20 students to win a new iPad. 3 students will be selected to win. What is the probablity that Manny, Moe and Jack will be the three selected?
Answer by checkley77(12844)

First pick=3/20
Second pick=2/19
Third pick=1/18
3/20*2/19*1/18=6/6,080=1/1,140 one chance in 1,140 that all 3 will win.